OpenStudy (anonymous):

Can anyone please help me? You need a 20% alcohol solution. On hand, you have a 400 mL of a 15% alcohol mixture. How much pure alcohol will you need to add to obtain the desired solution?

OpenStudy (anonymous):

you can set upthe equation as: x mL of 100% alchohol + 400 mL of 15% alcohol = (x + 400) mL of 20% alcohol

OpenStudy (anonymous):

you have 15% of 400 ml = \(.15\times 400=60\) ml of alcohol at present. if you add say \(x\) ml of alcohol you will have \(60+x\) ml of alcohol and \(400+x\) ml of solution you want \(400+x\) ml to be 20% alcohol, so set \[60+x=.20(400+x)\] and solve for \(x\)

OpenStudy (anonymous):

probably easiest to multiply by 10 on both sides to get \[600+10x=2(400+x)\] \[600+10x=800+2x\] \[8x=200\] etc

OpenStudy (anonymous):

THANK YOU VERY MUCH!

OpenStudy (anonymous):

yw hope steps are clear